OpenStudy (anonymous):

5/4(x - 1) I'm having trouble figuring out how to multiply the 5/4 and the -1

OpenStudy (anonymous):

it is simply \[\frac{ -5 }{ 4}\] or \[\frac{ -5 }{ 4 }x -\frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

Oh lol thanks

OpenStudy (anonymous):

What happens if I had to add 2

OpenStudy (anonymous):

give 2 a common denominator of 4 and add it to -5/4 so\[\frac{ -5 }{ 4 }+2=\frac{ -5 }{ 4 }+\frac{ 4 }{ 4 }=\frac{ -1 }{ 4 }\]

OpenStudy (anonymous):

No, the whole problem is added by 2

OpenStudy (anonymous):

5/4x - 5/4 + 2